. Given market demand Qd = 50 - P, and market supply P = Qs + 5


A) Find the market equilibrium price and quantity?


B) What would be the state of the market if market price was fixed at Birr 25 per unit?


C) Calculate and interpret price elasticity of demand at the equilibrium point.


1
Expert's answer
2022-01-27T02:02:34-0500

A. Qd= 50 - P, P = Qs + 5 -> Qs = P - 5. 

The market equilibrium price and quantity are:

Qd = Qs,

50 - P = P - 5,

2P = 55,

P = $27.5.

Q = 27.5 - 5 = 22.5 units.

B. If market price was fixed at $25 per unit, then there will be a shortage (Qd > Qs).

C. Price elasticity of demand at the equilibrium point is:

"Ed = -1×\frac{27.5}{22.5} = -1.22,"

so the demand is elastic.